Dent Corp. has just reported annual free-cash-flows of $1,500 million. Dent Corp. expects to have an investment ratio of 60% and an ROIC of 25% indefinitely. Dent Corp currently has $ 500 million in net debt, 300 million shares outstanding, and an annual weighted average cost of capital of 20.00%. Given this information, estimate the price of a share of Dent Corp.’s common stock using the discount free-cash-flow model. Round your final answer to two decimals.
Answer #1

Expected long term Growth rate (g) = investment ratio x ROIC

Expected long term Growth rate (g) = 60% x 25%

Expected long term Growth rate (g) = 15%

WACC = 20%

Annual free-cashflows (CF0) = $1500 million

Annual free-cashflows (CF1) = CF0 x (1+g)

Annual free-cashflows (CF1) = 1500 x (1+15%)

Annual free-cashflows (CF1) = $ 1725 million

Using Gordon's Dividend Growth model,

Value of firm = CF1 / (WACC-g)

Value of firm = 1725 / (20%-15%)

Value of firm = $ 34500 million

Market Value of Debt = $ 500 million

Value of Equity = Value of Firm - Value of Debt

Value of Equity = 34500-500

Value of Equity = $ 34,000 million

No of share outstanding = 300 million shares

Price of share = 34000 / 300 = $113.33

Thumbs up please if satisfied. Thanks :)

Comment if further doubts in above